# FROST-CLUSTERS – I. Hierarchical star cluster assembly boosts intermediate-mass black hole formation

### v1

@@ 1 Introduction
The so-called fast runaway collision scenario relies on a very massive star (VMS [^2] We refer any stars more massive than $m_{\star}>{150}\>\mathrm{M_{\odot}}$ as VMSs in this work.) growing via repeated mergers to sufficiently high masses (Lee, 1987; Quinlan & Shapiro, 1990; Bonnell et al., 1998) to produce an IMBH remnant above the (P)PISN mass gap range. The IMBH may further grow by accreting stellar material in tidal disruption events (TDEs) and by merging with other BHs. Numerical simulations of the fast scenario (e.g. Portegies Zwart et al. 1999; Portegies Zwart & McMillan 2002; Gürkan et al. 2004; Portegies Zwart et al. 2004; Freitag et al. 2006a, b; Rizzuto et al. 2021; Rizzuto et al. 2022; Arca Sedda et al. 2023a; Arca Sedda et al. 2023c, b) show that the stellar collisions can result in a formation of a VMS or a supermassive star ($m_{\star}>{10^{3}}\>\mathrm{M_{\odot}}$, e.g. Denissenkov & Hartwick 2014; Gieles et al. 2018) with a mass from hundreds to potentially tens of thousands of solar masses. In a true runaway collision sequence, each collision further increases the collision rate. Rapid growth is however possible even without fulfilling this criterion. Typically only 11112222 stars per cluster can grow via collisions (Baumgardt & Klessen, 2011). In the global instability variant of the fast runaway collision scenario, stellar collisions become relevant if the collision time-scale is shorter than the characteristic age of the system (Escala, 2021; Vergara et al., 2023).
@@
The fast and the slow collisional IMBH formation scenarios are not mutually exclusive. Even if the stellar collisions do not result in the formation of an IMBH via the fast VMS formation and collapse channel, repeated tidal disruption events or BH-BH collisions may still result in a formation of mass gap BHs or IMBHs (Morscher et al., 2015; Rodriguez et al., 2016; Askar et al., 2017; Banerjee, 2018; Fragione & Kocsis, 2018; Antonini et al., 2019; Rodriguez et al., 2019; Belczynski & Banerjee, 2020; Arca-Sedda et al., 2021; Banerjee, 2021; Rizzuto et al., 2021; Mapelli et al., 2021; Rizzuto et al., 2022; Rizzuto et al., 2023; Arca Sedda et al., 2023c). IMBH (or MBH seed) formation via the fast collisional runaway channel most likely occurs in young, massive, star clusters (Portegies Zwart et al., 2010) with low metallicity. First, low metallicity environments are favored due to weaker stellar winds of massive stars resulting in more massive IMBH remnants (e.g. Mapelli 2016; Di Carlo et al. 2021). There is still uncertainty about the wind mass-loss rates of massive stars. High wind mass-loss rates (especially at high metallicities) might suppress the VMS IMBH formation channel or inhibit it altogether (e.g. Belkus et al. 2007; Yungelson et al. 2008; Glebbeek et al. 2009; Pauldrach et al. 2012).
@@
Low-mass galaxies that host proto-globular clusters, observed with the aid of gravitational lensing (Livermore et al., 2017; Vanzella et al., 2019), present the densest star forming environments in the high-redshift Universe (Vanzella et al., 2023). The very recent James Webb Space Telescope (JWST) observations of the Cosmic Gems arc (Adamo et al., 2024) have revealed five massive star clusters at $z=10.2$ with masses of $M_{\star}\sim{10^{6}}\>\mathrm{M_{\odot}}$ and lensing-corrected sizes of $\sim 1$ pc, corresponding to high stellar surface densities of $\Sigma_{\star}\sim{10^{5}}\>\mathrm{M_{\odot}}\>\mathrm{pc}^{-2}$. The galaxy GN-z11 observed by the JWST at $z=10.6$ with super-solar N/O abundance at low metallicity (Bunker et al., 2023) might be an indication of early stages of globular cluster formation, or even a chemical fingerprint of supermassive stars (Charbonnel et al., 2023; Marques-Chaves et al., 2024).
The evolution of massive star clusters (Wang et al., 2016) and IMBH formation and growth within them (Rizzuto et al., 2023; Arca Sedda et al., 2023a) is typically studied in isolated simulation setups. There are a number of cosmological setups focusing on the formation of proto-globular clusters and their IMBH growth (Sakurai et al., 2017; Ma et al., 2020). However, the cosmological setups typically lack the necessary mass or spatial resolution to follow the collisional stellar dynamics at the level of the isolated setups.
@@
Observationally, IMBHs still remain elusive (Greene et al., 2020). Extrapolating the well-established the $M_{\bullet}$$\sigma$ relation (e.g. Ferrarese & Merritt 2000; Gebhardt et al. 2000) for massive galaxies into systems with lower masses and velocity dispersions, it has been proposed that dwarf galaxies, especially those with an active low-luminosity nucleus, might host IMBHs (Greene & Ho, 2007; Dong et al., 2012; Chilingarian et al., 2018; Greene et al., 2020; Mezcua et al., 2018; Reines, 2022). Several ultra- or hyper-luminous X-ray emitters (ULXs, HLXs) in nearby galaxies have been interpreted as candidates for accreting IMBHs with masses ranging from $m_{\bullet}\sim{200}\>\mathrm{M_{\odot}}$ to $m_{\bullet}\sim{10^{5}}\>\mathrm{M_{\odot}}$ (e.g. Colbert & Mushotzky 1999; Kaaret et al. 2001; Matsumoto et al. 2001; Strohmayer & Mushotzky 2003; Patruno et al. 2006; Farrell et al. 2009; Mezcua et al. 2013; Pasham et al. 2014; Mezcua et al. 2015; Mezcua 2017; Kim et al. 2020). ALMA observations of gas dynamics in dwarf elliptical NGC 404 (Davis et al., 2020) have revealed a BH with a mass of $m_{\bullet}\sim{5.5\times 10^{5}}\>\mathrm{M_{\odot}}$ consistent with both a high-mass IMBH or a low-mass MBH. In addition, there is stellar-dynamical evidence of dark objects in the centers of some globular clusters (GCs) and stripped dwarf galaxies in the IMBH mass range, consistent with concentrations of stellar-mass black holes, or possibly IMBHs (e.g. Gebhardt et al. 2002; Ibata et al. 2009; Noyola et al. 2010; van der Marel & Anderson 2010; Lützgendorf et al. 2011; Jalali et al. 2012; Lanzoni et al. 2013; Kamann et al. 2016; Baumgardt 2017; Pechetti et al. 2022; Della Croce et al. 2024). However, a number of studies do not find such stellar-dynamical evidence of IMBHs (Baumgardt et al., 2003; Baumgardt et al., 2019; Mann et al., 2019; Hénault-Brunet et al., 2020), and characteristic IMBH accretion signatures have not been observed in the clusters in question (Tremou et al., 2018). Still, the most definite evidence of IMBHs to date remains the gravitational-wave (GW) transient GW190521 (Abbott et al., 2020), a merger between two stellar-mass black holes with masses of $m_{\bullet}\sim{66}\>\mathrm{M_{\odot}}$ and $m_{\bullet}\sim{85}\>\mathrm{M_{\odot}}$ resulting in a formation of an IMBH with $m_{\bullet}\sim{150}\>\mathrm{M_{\odot}}$. In addition to gravitational waves, electromagnetic transients such as tidal disruption events may help to uncover the elusive IMBH population (e.g. Lin et al. 2018; Fragione & Kocsis 2018; Fragione et al. 2018).
@@ 2.1 Stellar collision rate: the cross section argument
[…]
@@ 2.4 Can the clusters retain their (IM)BHs?
only of the order of a few tens of km/s to $\sim 100$ km/s for star clusters in the mass range of $M_{\star}={10^{5}}\>\mathrm{M_{\odot}}$${10^{6}}\>\mathrm{M_{\odot}}$. Even Newtonian three-body and multiple interactions are occasionally strong enough to unbind BHs and stars from their host clusters, especially in low escape velocity environments of lower-mass star clusters and in clusters containing multiple IMBHs (Maliszewski et al. 2022, Souvaitzis et al., in prep.).
@@
We illustrate typical gravitational-wave recoil kick velocities compared to escape velocities from dense star clusters in Fig. 1. The recoil velocities are calculated using the fitting formulas by Zlochower & Lousto (2015). We use two spin direction options (aligned and anti-aligned spins), and three different secondary BH masses of $m_{\bullet}={10}\>\mathrm{M_{\odot}}$ (common low-mass BHs), $m_{\bullet}={60}\>\mathrm{M_{\odot}}$ (BHs just below the (P)PISN mass gap) and $m_{\bullet}={290}\>\mathrm{M_{\odot}}$ (IMBHs just above the (P)PISN mass gap). The primary BH mass $m_{\mathrm{p}}$ ranges in each model from $m_{\mathrm{s}}$ up to ${10^{4}}\>\mathrm{M_{\odot}}$ in the IMBH mass range. We assume the Geneva model for BH spins (Belczynski et al., 2020) as in their Eq. (3) with $Z=0.0004=0.02\>Z_{\odot}$.
From Fig. 1 it is evident that it is difficult to retain a merger remnant originating from a BH-BH binary that includes a BH below the mass gap ($m_{\mathrm{p}}\lesssim{63}\>\mathrm{M_{\odot}}$), unless the merging binary is equal-mass or nearly so, and has almost aligned spins. Primary IMBHs above the mass gap ($m_{\mathrm{p}}\gtrsim{290}\>\mathrm{M_{\odot}}$) can merge with low-mass BHs ($m_{\mathrm{s}}={10}\>\mathrm{M_{\odot}}$) with any relative spin orientation and be retained already in star clusters with masses $M_{\star}\sim{10^{3}}\>\mathrm{M_{\odot}}$${10^{4}}\>\mathrm{M_{\odot}}$. However, the primary IMBH masses above $m_{\bullet}\gtrsim{500}\>\mathrm{M_{\odot}}$ and $m_{\bullet}\gtrsim{1000}\>\mathrm{M_{\odot}}$ are required to retain the merger product in the star clusters with masses $M_{\star}={10^{6}}\>\mathrm{M_{\odot}}$ and $M_{\star}={10^{5}}\>\mathrm{M_{\odot}}$, respectively, after a merger with any stellar-mass BH below the mass gap. IMBH primaries with $m_{\mathrm{p}}\gtrsim{2000}\>\mathrm{M_{\odot}}$ are difficult to remove from any star cluster in mergers with stellar-mass BHs. Finally, IMBH-IMBH mergers may remove even massive primary IMBHs from their host clusters, provided that the IMBH formation channel results in multiple IMBHs in a single star cluster. A primary mass of $m_{\mathrm{p}}\sim{4500}\>\mathrm{M_{\odot}}$ is required to retain the remnant after a merger with an above-mass-gap secondary IMBH with $m_{\bullet}{290}\>\mathrm{M_{\odot}}$ in a star cluster with $M_{\star}={10^{5}}\>\mathrm{M_{\odot}}$.
@@ 3.1 Numerical methods: the updated BIFROST code
The main differences of the updated BIFROST code version used in this study compared to the original code of Rantala et al. (2023) are the following. First, a key element of the fourth-order forward integration, the calculation of the term involving the acceleration gradient, is simplified (appendix A.1). Next, the secular triple integration and the standard algorithmically regularized few-body integration (LogH) are replaced by an implementation of the slow-down algorithmic regularization (SDAR) integrator (appendix A.2). Finally, we present the coupling of the fast stellar population synthesis code SEVN (Iorio et al., 2023) into BIFROST allowing stellar evolution in the N-body simulations (appendix B.2). In this study, we focus on single stellar evolution in the initial stellar mass range of ${2.2}\>\mathrm{M_{\odot}}\lesssim m_{\star}\lesssim{600}\>\mathrm{M_{\odot}}$. Each of the three code updates is described in detail in the appendix.
@@ 3.2 Initial conditions: individual star cluster models
Thus, a size increase by a factor of up to $\sim 10$ is required from the embedded phase to the young massive cluster phase. The sizes of star clusters increase due to internal collisional dynamics, external tides and stellar evolution (e.g. Arca Sedda et al. 2023a), with gas expulsion suggested as an additional physical process contributing to the size growth (Banerjee & Kroupa, 2017).
@@
We illustrate the particle numbers and half-mass densities of a number of direct N-body (Wang et al., 2015; Di Carlo et al., 2021; Banerjee, 2021; Rastello et al., 2021; Rizzuto et al., 2021; Kamlah et al., 2022; Arca Sedda et al., 2023a) and cluster Monte Carlo (Askar et al., 2017; Rodriguez et al., 2016; Rodriguez et al., 2019; Kremer et al., 2020; Maliszewski et al., 2022) simulations following Fig. 1 of Arca Sedda et al. (2023a) in Fig. 2. We show both our isolated star cluster models, and the population of star clusters for the hierarchical cluster assembly simulations. The cluster assembly region simulations consist of individual star clusters with particle numbers in the range $1.7\times 10^{2}\lesssim N\lesssim 4.0\times 10^{5}$ particles, up to $N_{\mathrm{tot}}\sim 2.35\times 10^{6}$ particles per simulation. The isolated simulations we run have up to $N=1.0\times 10^{6}$ particles. Our adopted mass-size relation in Eq. (11) results in an initial star cluster population consistent with other recent simulation studies of massive star cluster evolution and IMBH formation.

### v2

@@ 1 Introduction
The so-called fast runaway collision scenario relies on a very massive star (VMS [^2] We refer any stars more massive than $m_{\star}>{150}\>\mathrm{M_{\odot}}$ as VMSs in this work.) growing via repeated mergers to sufficiently high masses (Lee, 1987; Quinlan & Shapiro, 1990; Bonnell et al., 1998) to produce an IMBH remnant above the (P)PISN mass gap range. The IMBH may further grow by accreting stellar material in tidal disruption events (TDEs) and by merging with other BHs. Numerical simulations of the fast scenario (e.g. Portegies Zwart et al. 1999; Portegies Zwart & McMillan 2002; Gürkan et al. 2004; Portegies Zwart et al. 2004; Freitag et al. 2006a, b; Rizzuto et al. 2021; Rizzuto et al. 2022; Arca Sedda et al. 2024a; Arca Sedda et al. 2023, 2024b; González Prieto et al. 2024) show that the stellar collisions can result in a formation of a VMS or a supermassive star ($m_{\star}>{10^{3}}\>\mathrm{M_{\odot}}$, e.g. Denissenkov & Hartwick 2014; Gieles et al. 2018) with a mass from hundreds to potentially tens of thousands of solar masses. In a true runaway collision sequence, each collision further increases the collision rate. Rapid growth is however possible even without fulfilling this criterion. Typically only 11112222 stars per cluster can grow via collisions (Baumgardt & Klessen, 2011). In the global instability variant of the fast runaway collision scenario, stellar collisions become relevant if the collision time-scale is shorter than the characteristic age of the system (Escala, 2021; Vergara et al., 2023).
@@
The fast and the slow collisional IMBH formation scenarios are not mutually exclusive. Even if the stellar collisions do not result in the formation of an IMBH via the fast VMS formation and collapse channel, repeated tidal disruption events or BH-BH collisions may still result in a formation of mass gap BHs or IMBHs (Morscher et al., 2015; Rodriguez et al., 2016; Askar et al., 2017; Banerjee, 2018; Fragione & Kocsis, 2018; Antonini et al., 2019; Rodriguez et al., 2019; Belczynski & Banerjee, 2020; Arca-Sedda et al., 2021b; Banerjee, 2021; Rizzuto et al., 2021; Mapelli et al., 2021; Rizzuto et al., 2022; Rizzuto et al., 2023; Arca Sedda et al., 2023). IMBH (or MBH seed) formation via the fast collisional runaway channel most likely occurs in young, massive, star clusters (Portegies Zwart et al., 2010) with low metallicity. First, low metallicity environments are favored due to weaker stellar winds of massive stars resulting in more massive IMBH remnants (e.g. Mapelli 2016; Di Carlo et al. 2021). There is still uncertainty about the wind mass-loss rates of massive stars. High wind mass-loss rates (especially at high metallicities) might suppress the VMS IMBH formation channel or inhibit it altogether (e.g. Belkus et al. 2007; Yungelson et al. 2008; Glebbeek et al. 2009; Pauldrach et al. 2012).
@@
Low-mass galaxies that host proto-globular clusters, observed with the aid of gravitational lensing (Livermore et al., 2017; Vanzella et al., 2019), present the densest star forming environments in the high-redshift Universe (Vanzella et al., 2023). The recent James Webb Space Telescope (JWST) observations have revealed clumpy, clustered star formation regions at redshifts beyond $z>6$, such as the Cosmic Grapes galaxy (Fujimoto et al., 2024), enriched by the young, massive stars (Topping et al., 2024). The JWST observations of the Cosmic Gems arc (Adamo et al., 2024) have revealed five massive star clusters at $z=10.2$ with masses of $M_{\star}\sim{10^{6}}\>\mathrm{M_{\odot}}$ and lensing-corrected sizes of $\sim 1$ pc, corresponding to high stellar surface densities of $\Sigma_{\star}\sim{10^{5}}\>\mathrm{M_{\odot}}\>\mathrm{pc}^{-2}$. The galaxy GN-z11 observed by the JWST at $z=10.6$ with super-solar N/O abundance at low metallicity (Bunker et al., 2023) might be an indication of early stages of globular cluster formation, or even a chemical fingerprint of supermassive stars (Charbonnel et al., 2023; Marques-Chaves et al., 2024).
The evolution of massive star clusters (Wang et al., 2016) and IMBH formation and growth within them (Rizzuto et al., 2023; Arca Sedda et al., 2024a) is typically studied in isolated simulation setups. There are a number of cosmological setups focusing on the formation of proto-globular clusters and their IMBH growth (Katz et al., 2015; Sakurai et al., 2017; Ma et al., 2020). However, the cosmological setups typically lack the necessary mass or spatial resolution to follow the collisional stellar dynamics at the level of the isolated setups.
@@
Observationally, IMBHs still remain elusive (Greene et al., 2020). Extrapolating the well-established the $M_{\bullet}$$\sigma$ relation (e.g. Ferrarese & Merritt 2000; Gebhardt et al. 2000) for massive galaxies into systems with lower masses and velocity dispersions, it has been proposed that dwarf galaxies, especially those with an active low-luminosity nucleus, might host IMBHs (Greene & Ho, 2007; Dong et al., 2012; Chilingarian et al., 2018; Greene et al., 2020; Mezcua et al., 2018; Reines, 2022). Several ultra- or hyper-luminous X-ray emitters (ULXs, HLXs) in nearby galaxies have been interpreted as candidates for accreting IMBHs with masses ranging from $m_{\bullet}\sim{200}\>\mathrm{M_{\odot}}$ to $m_{\bullet}\sim{10^{5}}\>\mathrm{M_{\odot}}$ (e.g. Colbert & Mushotzky 1999; Kaaret et al. 2001; Matsumoto et al. 2001; Strohmayer & Mushotzky 2003; Patruno et al. 2006; Farrell et al. 2009; Mezcua et al. 2013; Pasham et al. 2014; Mezcua et al. 2015; Mezcua 2017; Kim et al. 2020). ALMA observations of gas dynamics in dwarf elliptical NGC 404 (Davis et al., 2020) have revealed a BH with a mass of $m_{\bullet}\sim{5.5\times 10^{5}}\>\mathrm{M_{\odot}}$ consistent with both a high-mass IMBH or a low-mass MBH. In addition, there is stellar-dynamical evidence of dark objects in the centers of some globular clusters (GCs) and stripped dwarf galaxies in the IMBH mass range, consistent with concentrations of stellar-mass black holes, or possibly IMBHs (e.g. Gebhardt et al. 2002; Ibata et al. 2009; Noyola et al. 2010; van der Marel & Anderson 2010; Lützgendorf et al. 2011; Jalali et al. 2012; Lanzoni et al. 2013; Kamann et al. 2016; Baumgardt 2017; Pechetti et al. 2022; Della Croce et al. 2024). However, a number of studies do not find such stellar-dynamical evidence of IMBHs (Baumgardt et al., 2003; Baumgardt et al., 2019; Mann et al., 2019; Hénault-Brunet et al., 2020), and characteristic IMBH accretion signatures have not been observed in the clusters in question (Tremou et al., 2018). Still, the most definite evidence of IMBHs to date remains the gravitational-wave (GW) transient GW190521 (Abbott et al., 2020), a merger between two stellar-mass black holes with masses of $m_{\bullet}\sim{66}\>\mathrm{M_{\odot}}$ and $m_{\bullet}\sim{85}\>\mathrm{M_{\odot}}$ resulting in a formation of an IMBH with $m_{\bullet}\sim{150}\>\mathrm{M_{\odot}}$. In addition to gravitational waves, electromagnetic transients such as tidal disruption events may help to uncover the elusive IMBH population (e.g. Lin et al. 2018; Fragione & Kocsis 2018; Fragione et al. 2018b).
@@ 2.1 Stellar collision rate: the cross section argument
Instead, gravitational focusing is important and the collision rate is inversely proportional to $\sigma$, and the collision rate decreases in high velocity dispersion environments. The collision rates predicted by the simple cross section argument of Eq. (1) are typically low. For example, a dense star cluster following a Plummer (1911) density profile with a total mass of $M_{\star}={2\times 10^{5}}\>\mathrm{M_{\odot}}$ and half-mass radius of $r_{\mathrm{h}}=0.25$ pc consisting of equal-mass main-sequence stars of $m={10}\>\mathrm{M_{\odot}}$ has a collision rate of only $t_{\mathrm{coll}}^{\mathrm{-1}}\sim 10^{-3}$ Myr-1.
@@ 2.4 Can the clusters retain their (IM)BHs?
only of the order of a few tens of km/s to $\sim 100$ km/s for star clusters in the mass range of $M_{\star}={10^{5}}\>\mathrm{M_{\odot}}$${10^{6}}\>\mathrm{M_{\odot}}$. Even Newtonian three-body and multiple interactions are occasionally strong enough to unbind BHs and stars from their host clusters, especially in low escape velocity environments of lower-mass star clusters and in clusters containing multiple IMBHs (Holley-Bockelmann et al. 2008; Fragione et al. 2018a; Maliszewski et al. 2022; Arca Sedda et al. 2023, Souvaitzis et al., in prep.).
@@
We illustrate typical gravitational-wave recoil kick velocities compared to escape velocities from dense star clusters in Fig. 1. The recoil velocities are calculated using the fitting formulas provided by Zlochower & Lousto (2015). We use the same formulas in BIFROST as briefly described in Rantala et al. (2023). We use two spin direction options (aligned and anti-aligned spins), and three different secondary BH masses of $m_{\bullet}={10}\>\mathrm{M_{\odot}}$ (common low-mass BHs), $m_{\bullet}={60}\>\mathrm{M_{\odot}}$ (BHs just below the (P)PISN mass gap) and $m_{\bullet}={290}\>\mathrm{M_{\odot}}$ (IMBHs just above the (P)PISN mass gap). The primary BH mass $m_{\mathrm{p}}$ ranges in each model from $m_{\mathrm{s}}$ up to ${10^{4}}\>\mathrm{M_{\odot}}$ in the IMBH mass range. As in BIFROST, we assume the Geneva model for BH spins (Belczynski et al., 2020) as in their Eq. (3) with $Z=0.0004=0.02\>Z_{\odot}$.
From Fig. 1 it is evident that it is difficult to retain a merger remnant originating from a BH-BH binary that includes a BH below the mass gap ($m_{\mathrm{p}}\lesssim{63}\>\mathrm{M_{\odot}}$), unless the merging binary is equal-mass or nearly so, and has almost aligned spins. Primary IMBHs above the mass gap ($m_{\mathrm{p}}\gtrsim{290}\>\mathrm{M_{\odot}}$) can merge with low-mass BHs ($m_{\mathrm{s}}={10}\>\mathrm{M_{\odot}}$) with any relative spin orientation and be retained already in star clusters with masses $M_{\star}\sim{10^{3}}\>\mathrm{M_{\odot}}$${10^{4}}\>\mathrm{M_{\odot}}$. However, the primary IMBH masses above $m_{\bullet}\gtrsim{500}\>\mathrm{M_{\odot}}$ and $m_{\bullet}\gtrsim{1000}\>\mathrm{M_{\odot}}$ are required to retain the merger product in the star clusters with masses $M_{\star}={10^{6}}\>\mathrm{M_{\odot}}$ and $M_{\star}={10^{5}}\>\mathrm{M_{\odot}}$, respectively, after a merger with any stellar-mass BH below the mass gap. IMBH primaries with $m_{\mathrm{p}}\gtrsim{2000}\>\mathrm{M_{\odot}}$ are difficult to remove from any star cluster in mergers with stellar-mass BHs. For further discussion about merging IMBHs with stellar mass BHs see e.g. Arca Sedda et al. (2021a). Finally, IMBH-IMBH mergers may remove even massive primary IMBHs from their host clusters, provided that the IMBH formation channel results in multiple IMBHs in a single star cluster. A primary mass of $m_{\mathrm{p}}\sim{4500}\>\mathrm{M_{\odot}}$ is required to retain the remnant after a merger with an above-mass-gap secondary IMBH with $m_{\bullet}={290}\>\mathrm{M_{\odot}}$ in a star cluster with $M_{\star}={10^{5}}\>\mathrm{M_{\odot}}$.
@@ 3.1 Numerical methods: the updated BIFROST code
The main differences of the updated BIFROST code version used in this study compared to the original code of Rantala et al. (2023) are the following. First, a key element of the fourth-order forward integration, the calculation of the term involving the acceleration gradient, is simplified (appendix A.1). Next, the secular triple integration and the standard algorithmically regularized few-body integration (LogH) are replaced by an implementation of the slow-down algorithmic regularization (SDAR) integrator (appendix A.2). Finally, we present the coupling of the fast stellar population synthesis code SEVN (Iorio et al., 2023) into BIFROST allowing stellar evolution in the N-body simulations (appendix B.2). In this study, we focus on single stellar evolution in the initial stellar mass range of ${2.2}\>\mathrm{M_{\odot}}\lesssim m_{\star}\lesssim{600}\>\mathrm{M_{\odot}}$. Stars in dynamically formed binaries evolve as single stars, ignoring any binary stellar evolution physics. Setups with primordial binary stars and full binary stellar evolution will be explored in a forthcoming study. Each of the three code updates is described in detail in the appendix.
@@ 3.2 Initial conditions: individual star cluster models
Thus, a size increase by a factor of up to $\sim 10$ is required from the embedded phase to the young massive cluster phase. The sizes of star clusters increase due to internal collisional dynamics, external tides and stellar evolution (e.g. Arca Sedda et al. 2024a), with gas expulsion suggested as an additional physical process contributing to the size growth (Banerjee & Kroupa, 2017).
@@
We illustrate the particle numbers and half-mass densities of a number of direct N-body (Wang et al., 2015; Di Carlo et al., 2021; Banerjee, 2021; Rastello et al., 2021; Rizzuto et al., 2021; Kamlah et al., 2022; Arca Sedda et al., 2024a) and cluster Monte Carlo (Askar et al., 2017; Rodriguez et al., 2016; Rodriguez et al., 2019; Kremer et al., 2020; Maliszewski et al., 2022) simulations following Fig. 1 of Arca Sedda et al. (2024a) in Fig. 2. We show both our isolated star cluster models, and the population of star clusters for the hierarchical cluster assembly simulations. The cluster assembly region simulations consist of individual star clusters with particle numbers in the range $1.7\times 10^{2}\lesssim N\lesssim 4.0\times 10^{5}$ particles, up to $N_{\mathrm{tot}}\sim 2.35\times 10^{6}$ particles per simulation. The isolated simulations we run have up to $N=1.0\times 10^{6}$ particles. Our adopted mass-size relation in Eq. (11) results in an initial star cluster population consistent with other recent simulation studies of massive star cluster evolution and IMBH formation.